# Unravelling the Mechanisms of Manipulating Numbers in Language Models

### v1

@@ Abstract
Recent work has shown that different large language models (LLMs) converge to similar and accurate input embedding representations for numbers. These findings conflict with the documented propensity of LLMs to produce erroneous outputs when dealing with numeric information. In this work, we aim to explain this conflict by exploring how language models manipulate numbers and quantify the lower bounds of accuracy of these mechanisms. We find that despite surfacing errors, different language models learn interchangeable representations of numbers that are systematic, highly accurate and universal across their hidden states and the types of input contexts. This allows us to create universal probes for each LLM and to trace information — including the causes of output errors — to specific layers. Our results lay a fundamental understanding of how pre-trained LLMs manipulate numbers and outline a potential of more accurate probing techniques in addressed refinements of LLMs’ architectures.
@@ 1 Introduction
One major limitation of large language models (LLMs) is their inability to guarantee factually precise and correct outputs. This is a fundamental trait due to the probabilistic nature of transformer models, and thus not a solvable problem, as has recently been emphasized by (Kalai et al., 2025). However, knowing the extent of the problem (and whether it can be mitigated) is a feasible task; one that is important given the recent ubiquity of LLMs in research and broader society.
Mathematical problems provide an ideal playground in which to assess the current ability of LLMs to manipulate precise information. Because it operates through a set of formal rules and produces computationally verifiable outputs, mathematics provides the opportunity to rate not only the tendency of a model to output incoherent information (Sun et al., 2024), but also to examine in detail the coherence of the reasoning steps outputted by a given LLM. While this research direction is rife with challenges, ranging from manipulating implicitly visual data in geometry problems (Ahn et al., 2024) to mapping plain language onto specific types of problems (Duchnowski et al., 2025), it has also been met by genuine successes, e.g., as an aid to solve mathematical problems (Li et al., 2025).
A large body of work has highlighted limitations of LLMs when it comes to arithmetic tasks (Xu et al., 2025; Kvinge et al., 2025; Satpute et al., 2024; Lee et al., 2025; Bertolazzi et al., 2025; Nikankin et al., 2025), which in turn has led to various proposals to alleviate these limitations centered on re-designing numeric representations to be principled and precise (e.g., Charton, 2022; Feng et al., 2024; Golkar et al., 2023). More recent developments have suggested that the issues models encounter with arithmetic need not be tied to the quality of their representations of numbers. Several works have attempted to characterize the types of representations models rely on to perform arithmetic computations (Wennberg & Henter, 2024; Zhou et al., 2024a; Zhu et al., 2025; Kadlčík et al., 2025). This research direction has proposed that number representations stand out in terms of their structure, with Zhou et al. (2024a) and Kadlčík et al. (2025) arguing that models rely on sinusoidal structures, whereas Zhu et al. (2025) suggest that representations are better described using a linear model than a multi-layer perceptron. These results have also been combined with observations as to how models process and succeed at arithmetic tasks.
In this paper, we provide evidence that LLMs converge to systematic and accurate sinusoidal representations that are maintained throughout the model in different forms. These representations are largely equivalent across LLMs of different sizes and families. Building upon this, we find that numeric information from LLMs can be accurately extracted using universal probes and show that more accurate numeric probes allow us to attribute up to 94% of model errors in arithmetical reasoning to particular layers, overriding the existing, correct results. [^1] We make all our methods and analyses available to any use at the project repository: https://github.com/prompteus/numllama
@@ 2 Related work
Math capabilities of LLMs
The seminal remarks of Brown et al. (2020) that LLMs that are large enough develop arithmetic capabilities have ushered in a large body of work aimed at evaluating the mathematical capabilities of LLMs (Hendrycks et al., 2021; Cobbe et al., 2021; Sun et al., 2024; Yu et al., 2024) or lack thereof (Nikankin et al., 2025). This in turn has encouraged a focus on how to bolster the capabilities of LLMs or interpret how LLMs perform computations (Zhang et al., 2024; Stolfo et al., 2023). Most relevant to ours, some of the previous work especially focuses on how numbers should be represented in principle. For instance, Charton (2022) assess an impact on models performance in linear algebra problems when employing different encoding schemes based on scientific notation on linear algebra problems. Golkar et al. (2023) propose to encode numeric values by incorporating a scaled, learned control token <NUM>. Feng et al. (2024) remark that the precision of the numeric representation type impacts performance on arithmetic tasks, and argue to use the logarithmic-precision architecture of Feng et al. (2023).
@@
The increased focus on evaluating LLM performance in verifiable domains such as math problems has also led to increased interest in the mechanisms by which LLMs compute arithmetic functions. Nanda et al. (2023) demonstrate how a model trained from scratch on modular addition relies on trigonometric operation: it maps inputs onto a unit circle, corresponding to a specific rotation, and then learns to combine the two rotations to derive a valid solution. Kantamneni & Tegmark (2025) apply circuit analysis to highlight how general pretrained models perform addition, and highlight that representations are mapped onto a helix that can be manipulated using trigonometric operations. Zhou et al. (2024b) recast these observations in terms of Fourier components (since Fourier transformations map arbitrary functions unto combinations of periodic trigonometric functions), and highlight the distinct role of attention and feedforward sublayers. Focusing more narrowly on representations rather than processing, Zhu et al. (2025) argue that non-linear probes do not provide a better fit than linear probes. Levy & Geva (2025) remark that it is possible to retrieve the digit (value mod 10) of numeric inputs with high accuracy. Kadlčík et al. (2025) propose a linear probe architecture that factors in the sinusoidal nature of the embeddings of numbers, and show that they can retrieve numeric precision with high accuracy.
In summary, there is a growing body of evidence stating that models rely on trigonometric representations for numbers. Here, we build upon Kadlčík et al. (2025) specifically, as the high accuracy of their proposed probe provides new opportunities and novel research directions — as we find, also in tracking and verifying the accuracy of a model’s inner representations throughout its computations, robust across contexts, model types and operations.
@@ 3 Universality of Sinusoidal Representations of Numbers
3.1 Models Learn Equivalent Representations of Numbers
The first point we address is the distinctiveness of LLMs’ representations of numbers. Results from Kadlčík et al. (2025) and Zhu et al. (2025) suggest that different models converge to the same type of number representations, but they do not provide a direct assessment of how closely the representations match across models – which could evidence that a shared representation is a causal consequence of architectural bias and optimization process rather than a coincidental artifact. Following Kadlčík et al., we focus on the input embeddings from eight LLMs of diverse sizes and families with open-sourced, pre-trained checkpoints: OLMo 2 (OLMo et al., 2025), Llama 3 (Grattafiori et al., 2024) and Phi 4 (Abdin et al., 2024).
@@ 3.1 Models Learn Equivalent Representations of Numbers
Figure 1: Representational similarity analysis (RSA) scores
To quantify whether models converge to similar embeddings, we start by conducting a simple Representational Similarity Analysis (RSA; Kriegeskorte et al., 2008) across the input embeddings for number word-pieces in paired models. [^2] RSA is a second-order similarity measurement. Let $A=\{\mathbf{a}_{1},\dots\mathbf{a}_{n}\}$ and $B=\{\mathbf{b}_{1},\dots\mathbf{b}_{n}\}$ be two sets of matching representations, in our case the input embeddings for number word-pieces in two models $\theta_{A}$ and $\theta_{B}$. We can assess the extent to which $A$ and $B$ encode the same type of similarity structures by simply computing the (Spearman) correlation $\rho_{\mathbf{s}_{A},\mathbf{s}_{B}}$, where $\mathbf{s}_{A}=\left(\cos(\mathbf{a}_{1},\mathbf{a}_{2})\dots\cos(\mathbf{a}_{n-1},\mathbf{a}_{n})\right)$ and $\mathbf{s}_{B}=\left(\cos(\mathbf{b}_{1},\mathbf{b}_{2})\dots\cos(\mathbf{b}_{n-1},\mathbf{b}_{n})\right)$ are two vectors tracking all pairwise similarities within $A$ and $B$ respectively. As a baseline, we compute RSA scores for a random sample of 1000 pieces present in the vocabularies of all the models under consideration. Results of the analysis are displayed in Figure 1. Across all pairs of models, we find that number embeddings systematically yield higher RSA scores than a random sample of word pieces. While some models (especially in the Phi family) are not as well aligned with other models, we do observe high scores both within and across families, demonstrating that models converge to embedding spaces with equivalent similarity structure.
@@
Another approach to quantifying the similarity of frequential representations characteristic for numbers Kadlčík et al. (2025); Zhou et al. (2024a) is by inspecting their base frequencies through Fourier decompositions. Applying a PCA transformation followed by a Fourier transform allows us to quantify the magnitude of individual frequencies (LABEL:fig:fft_iou:example). Subsequently, we determine whether two models agree as to how they rank frequencies by computing a simple intersection-over-union (IoU) of the top $k$ frequencies. We repeat this process for every pair of models, using the input embeddings of number pieces (LABEL:fig:fft_iou:num) as well as that of random pieces as previously (LABEL:fig:fft_iou:rand). When considering the top $k=63$ frequencies, [^3] See Section A.1 for details how this value of $k$ is found. we find perfect agreement across all models in terms of number pieces.
In summary, our analyses show that the LLMs from different families learn to represent numbers in a similar topology and trigonometric features using the same dominant frequencies. This is specific to numbers – we do not reproduce these observations with other word pieces shared across models.
@@ 3.2 Number Representations are Always Sinusoidal
In subsection 3.1, we presented evidence that the input embeddings of numbers in LLMs have certain properties related to how they encode numeric information, such as being sinusoidal. In this section, we aim to find whether such representations are also maintained and employed throughout the models.
@@
For qualitative assessment, we visualize the PCA and its Fourier transform of internal activations of numerical tokens in Section A.3. Besides the apparent wave-like pattern, the representations have a sparse Fourier transform, confirming the sinusoidal character. For quantitative analysis, we look at the model’s internal representations through the lens of the sinusoidal probe proposed by Kadlčík et al. (2025). This probe was designed to map input embeddings of LLMs into an integer, thus classifying embeddings into predefined range of numeric values. The sinusoidal probe is defined as:
$f_{\sin}(\mathbf{x})$ $=(\mathbf{W}_{\mathrm{out}}\mathbf{S})^{T}(\mathbf{W}_{\mathrm{in}}\mathbf{x})$ (1)
$\mathbf{S}_{ij}$ $=\begin{cases}\sin(ie^{j}1000/d)\qquad\mathrm{if~}j\equiv 0\mod 2\\ \cos(ie^{j+1}1000/d)\quad\mathrm{if~}j\equiv 1\mod 2\end{cases}$
where $\mathbf{W}_{\mathrm{in}}$ and $\mathbf{W}_{\mathrm{out}}$ are learned parameters, and $\mathbf{S}$ injects an inductive bias in the classifier towards sinusoidal representations. We build upon an assumption verified by Kadlčík et al. (2025) stating that sinusoidal probes indeed adapt and employ a sinusoidal representation from inductive bias, if a sinusoidal representation is also present in the input embeddings.
@@
We first assess whether sinusoidal probes are the most suitable, i.e., an accurate choice for decoding the internal representations of the model. To contextualize the accuracy in terms of sinusoidal quality of the representation, we also evaluate other types of probes used to decode numbers in previous work Feng et al. (2023); Zhu et al. (2025); Kadlčík et al. (2025) as baselines. we train the sinusoidal probe on hidden representations of each layer of Llama 3.2 1B model when processing addition prompts of the form $x_{1}$ + $x_{2}$’, where $x_{1}$ and $x_{2}$ are integers. We specifically evaluate whether the value of $x_{2}$ can be retrieved from the model’s hidden representation on each layer. To assess generalization rather than memorization capacity, we split the prompts and corresponding representations into training, validation and test sets containing distinct extracted numbers ($x_{2}$).
LABEL:fig:per_layer_acc_across_probes shows the accuracy of different probes extracting the value of input numbers from representations of each model’s layer, evaluated across six different models. The superior accuracy of sinusoidal probe provides an indicator of both (i) the sinusoidal character, and (ii) accuracy of input number representation across layers.
@@
Figure 3: Accuracy of decoding numeric input token from internal activations of language models
Thus far, an important caveat of ours, as well as a methodology of previous work probing representations of numbers in previous work (section 2) is that we focus on a very specific type of input, namely additions, without any sort of natural language to contextualize them. As such, it is reasonable to assume that the results presented in Figure 3 may not carry to a more natural context: there is little guarantee that the behavior of a model on arithmetic problems is indicative of its behavior a broad range of applications requiring manipulations of numbers in natural language contexts.
To evaluate the impact of natural language context on accuracy and regularity of numeric representations, we curate a dataset covering four domains involving intensive numeric manipulations: arithmetic (mathematical word problems), temporal (date extraction), medical (ICD-10 diagnostic codes) and culinary (recipe ingredients/quantities). The complete dataset specifications are detailed in Table 1 in Appendix B. To maximize the quality of learned probes and thus, representativeness of our findings, we also tackle the natural-language class imbalance; as not every number has equal probability to occur in a date or in a recipe, we replace original numbers with integers uniformly sampled from the model vocabulary. We then fit separate probes for activations from every layer, again holding out a set of 100 randomly selected numbers for evaluating generalization.
Fig. LABEL:fig:per_layer_acc_across_models_natural displays the accuracies of input value extraction from model’s hidden states within natural-language contexts across six LLMs of different sizes and three different families. As we can see, the sinusoidal probes reach over 70% of accuracy in all but three cases and over 90% of accuracy for a majority of probing scenarios. Taking into an account that the learned sinusoidal probe also induce a certain error. Therefore, these accuracies present a lower-bound of accuracy of models’ representation of input numbers in natural contexts in a sinusoidal representation — showing that the accurate, sinusoidal representation is indeed characteristic and employed by a wide range of recent language models.
A vast majority of cases performing lower than 80% occurs in probing the models’ first and a last layer — we analyse these cases in detail later in Section 3.3.
@@
Figure 4: Generalization of probes fitted on natural-language occurrences of numeric tokens (solid line), and synthetic, mathematical contexts (dashed line).
@@
A comparison of probes’ performance in math an natural-language contexts (Figure 3) may suggest that a type of context does not influence the results of probing accuracy. However, in Figure 4, we show that probes fitted with natural-language contexts are much more robust in application across both math and natural-language contexts; While both the math and natural-language probes generalize well within their training context type, the natural-language probes are applicable comparably even under a substantial context distribution shift, suggesting that a wider variety of contexts serves as an effective regularization strategy. This observation draws two important implications: First, it informs future work tracing models’ internal computations to favor probes trained in natural-language settings. Second, it restrains future work in interpretability from drawing broad conclusions on models’ mechanics from fully-synthetic settings, evidencing that such conclusions may not generalize to real-world applications.
@@
In short, our results underscore that the same type of sinusoidal representation of numbers holds in general across different model types. Numbers are represented in a similar, systematic fashion, regardless of which layer or type of contexts we consider.

### v2

# Language Models Learn Universal Representations of Numbers and Here’s Why You Should Care

@@ Abstract
Prior work has shown that large language models (LLMs) often converge to accurate input embedding for numbers, based on sinusoidal representations. In this work, we quantify that these representations are in fact strikingly systematic, to the point of being almost perfectly universal: different LLM families develop equivalent sinusoidal structures, and number representations are broadly interchangeable in a large swathe of experimental setups. We show that properly factoring in this characteristic is crucial when it comes to assessing how accurately LLMs encode numeric and other ordinal information, and that mechanistically enhancing this sinusoidality can also lead to reductions of LLMs’ arithmetic errors.
Language Models Learn Universal Representations of Numbers
and Here’s Why You Should Care
Michal Štefánik∗ 1,2,3 Timothee Mickus∗ 2 Marek Kadlčík∗ 3 Bertram Højer4 Michal Spiegel3,5 Raúl Vázquez2 Aman Sinha6 Josef Kuchař3 Philipp Mondorf7,8 Pontus Stenetorp1,9 1R&D Centre for Large Language Models, National Institute of Informatics, Japan 2University of Helsinki 3TransformersClub @ Faculty of Informatics, Masaryk University 4IT University of Copenhagen 5Kempelen Instutite of Information Technology 6IECL-ATILF, Université de Lorraine - ICANS Strasbourg 7MaiNLP, Center for Information and Language Processing, LMU Munich 8Munich Center for Machine Learning (MCML) 9University College London Equal contributors
@@ 1 Introduction
Mathematical problems provide an ideal playground in which to assess the current ability of large language models (LLMs) to manipulate precise information. While attempts to foster and improve mathematical reasoning in LLMs have been met with genuine empirical successes, e.g., as an aid to solving mathematical problems (Li et al., 2025), this field is still rife with challenges, particularly in ensuring that LLMs tackling mathematical problems remain robust and reliable. This calls for a deeper understanding of how robust the internal mechanisms of LLMs really are in operating over mathematical information.
Towards this goal, prior work identified that number representations stand out in terms of their sinusoidal structure (Nanda et al., 2023; Zhou et al., 2024; Levy and Geva, 2025; kadlčík2025pretrainedlanguagemodelslearn), accuracy of input embeddings (kadlčík2025pretrainedlanguagemodelslearn), and systematicity of internal mechanisms in addition (Kantamneni and Tegmark, 2025; Zhou et al., 2024). While these works have provided pivotal insights as to how LLMs represent and manipulate numbers, their scope is limited: kadlčík2025pretrainedlanguagemodelslearn’s findings only pertain to input embeddings; Zhou et al. find similarity of representations across models but do not quantify their alignment; Levy and Geva focus only on decimals as one particular aspect of representations. Crucially, no prior work considers natural language contexts, which strongly curtails the generalizability and practical usefulness of their findings, needed for the development of tools able to interpret – and verify – LLMs in practice.
To address these gaps and create a holistic picture of how language models handle numbers, this paper presents two main findings. First, we quantify that sinusoidal structures are almost perfectly universal across LLMs: models of different sizes and families converge to sinusoidal representations that are universal across models (§3.1) and layers (§3.3) while remaining highly systematic and accurate throughout the model’s internal processing (§3.2) – showing it is possible to track the processing of numeric information throughout models.
Second, we demonstrate the necessity of suitable explainability tools to make sense of how LLMs process numbers: we show how accounting for the sinusoidal nature of representations allows us to unravel the mechanism of manipulating large, multi-token numbers (§4.1) as well as other ordinal data (§4.3) in LLMs, and we mechanistically prove that a precision of sinusoidality in the model’s internal representation determines its output accuracy across arithmetic tasks (§4.2).
In summary, our main contributions are:
  1. 1.
    We develop a new, more accurate probe of numeric values that generalizes well across different tasks and models
  2. 2.
    We quantify that numeric representations of numbers are sinusoidal and highly accurate throughout the model, beyond the input embedding representations;
  3. 3.
    We are the first to analyze the LLMs’ numeric processing in natural-language settings and show that training numeric probes with real-world data is crucial for creating generalizing, practically useful probes;
  4. 4.
    We investigate mechanisms of numeric manipulations beyond addition and provide a mechanistic evidence that sinusoidality is crucial for models’ accuracy.
We make the new probe, with all our methods and analyses publicly available for any use. [^1] https://github.com/prompteus/numllama
@@ 2 Related work
Mathematical capabilities of LLMs
The seminal findings of Brown et al. (2020) on remarkable arithmetical abilities of LLMs have ushered in a large body of work aimed at evaluating more complex mathematical capabilities of LLMs (Hendrycks et al., 2021; Cobbe et al., 2021; Sun et al., 2024; Yu et al., 2024) or lack thereof (Nikankin et al., 2025). This in turn has encouraged a focus on how to bolster the capabilities of LLMs or interpret how LLMs perform computations (Zhang et al., 2024; Stolfo et al., 2023). Most relevant to ours, some of the previous work focuses specifically on how numbers should be represented in principle. For instance, Charton (2022) assesses an impact on models performance in linear algebra problems when employing different encoding schemes based on scientific notation on linear algebra problems. Golkar et al. (2023) propose to encode numeric values by incorporating a scaled, learned control token <NUM>. Feng et al. (2024) remark that the precision of the numeric representation type impacts performance on arithmetic tasks, and argue to use the logarithmic-precision architecture of Feng et al. (2023).
@@
The increased focus on evaluating LLM performance in verifiable domains such as math problems has also led to increased interest in the mechanisms by which LLMs compute arithmetic functions. Nanda et al. (2023) demonstrate how a model trained from scratch on modular addition relies on trigonometric operation: it maps inputs onto a unit circle, corresponding to a specific rotation, and then learns to combine the two rotations to derive a valid solution. Kantamneni and Tegmark (2025) apply circuit analysis to understand how general pretrained models perform addition, and report that representations are mapped onto a helix that can be manipulated using trigonometric operations. Zhou et al. (2024) refine these observations in qualitative analysis of Fourier components in LLMs’ representations (note that Fourier transformations map arbitrary functions into combinations of periodic trigonometric functions), and highlight the distinct role of attention and feedforward sublayers. Focusing more narrowly on representations rather than processing, Zhu et al. (2025) argue that non-linear probes do not provide a better fit than linear probes. Levy and Geva (2025) remark that it is possible to retrieve the digit (value mod 10) of numeric inputs with high accuracy. Our work builds upon kadlčík2025pretrainedlanguagemodelslearn that propose a probe architecture that factors in the sinusoidal nature of representations, and show that they can retrieve numeric values from input embeddings with much higher accuracy.
@@ 3 Universality of Sinusoidal Representations of Numbers
3.1 Different Models Learn Equivalent Representations of Numbers
The first point we address is the distinctiveness of LLMs’ representations of numbers. Results from kadlčík2025pretrainedlanguagemodelslearn and Zhu et al. (2025) suggest that different models converge to the same type of number representations, but they do not provide a direct assessment of how closely the representations match across models – which could evidence that a shared representation is a causal consequence of architectural bias and optimization process rather than a coincidental artifact. Following kadlčík2025pretrainedlanguagemodelslearn, we focus on the input embeddings from eight LLMs of diverse sizes and families with open-sourced, pre-trained checkpoints: OLMo 2 (OLMo et al., 2025), Llama 3 (Grattafiori et al., 2024) and Phi 4 (Abdin et al., 2024).
@@ 3.1 Different Models Learn Equivalent Representations of Numbers
Figure 1: RSA scores for input embeddings of LLMs
To quantify whether models converge to similar embeddings, we start by conducting a simple Representational Similarity Analysis (RSA; Kriegeskorte et al., 2008) across the input embeddings for number word-pieces in paired models. As a baseline, we compute RSA scores for a random sample of 1000 pieces present in the vocabularies of all the models under consideration. Results of the analysis are displayed in Figure˜1. Across all pairs of models, we find that number embeddings systematically yield consistently higher RSA scores than a random sample of word pieces. While some models (especially in the Phi family) are not as well aligned with other models, we do observe high scores both within and across families, demonstrating that models converge to embedding spaces with equivalent similarity structure.
@@
Another approach to quantifying the similarity of number representations is to inspect their base frequencies through Fourier decompositions (kadlčík2025pretrainedlanguagemodelslearn; Zhou et al., 2024). Applying a PCA transformation followed by a Fourier transform allows us to quantify the magnitude of individual frequencies (LABEL:fig:fft_iou:example). Subsequently, we determine whether two models agree as to how they rank frequencies by computing a simple intersection-over-union (IoU) of the top $k$ frequencies. We repeat this process for every pair of models, using the input embeddings of number pieces (LABEL:fig:fft_iou:num) as well as that of random pieces (LABEL:fig:fft_iou:rand). When considering the top $k=63$ frequencies, we find perfect agreement across all models for numbers. This provides a quantitative basis for some of Zhou et al.’s (2024; §C.1) qualitative observations. [^2] See our Section A.1 for technical details, including how the value of $k$ is set.
In summary, our analyses show that the LLMs from different families learn to represent numbers in the same topology and trigonometric features using the same dominant frequencies. This is specific to numbers: we do not reproduce these observations with other word pieces shared across models.
@@ 3.2 Number Representations are Always Sinusoidal
In Section˜3.1, we presented evidence that the input embeddings of numbers across different LLMs are highly consistent. We now focus on whether the sinusoidal mechanism of representations is maintained and employed throughout the model.
@@
For qualitative assessment, we visualize the PCA and its Fourier transform of internal activations of numerical tokens in Section˜A.4. Besides the apparent wave-like pattern, the representations have a sparse Fourier transform, confirming the sinusoidal character. For quantitative analysis, we look at the model’s internal representations through the lens of the sinusoidal probe proposed by kadlčík2025pretrainedlanguagemodelslearn. This probe was designed to map input embeddings of LLMs into an integer, thus classifying embeddings into predefined range of numeric values. It is defined as:
$\begin{aligned} f_{\sin}(\mathbf{x})&=(\mathbf{W}_{\mathrm{out}}\mathbf{S})^{T}(\mathbf{W}_{\mathrm{in}}\mathbf{x})\\ \mathbf{S}_{ij}&=\begin{cases}\sin(ie^{j}1000/h)\hfill\mathrm{if~}j\equiv 0\mod 2\\ \cos(ie^{j+1}1000/h)~\mathrm{if~}j\equiv 1\mod 2\end{cases}\end{aligned}$
where $\mathbf{W}_{\mathrm{in}}:h\times d$ and $\mathbf{W}_{\mathrm{out}}:h\times d$ are learned parameters, and $\mathbf{S}:h\times 1000$ injects an inductive bias in the classifier towards sinusoidality. [^3] Unless otherwise stated, we use $d=100$; $h$ corresponds to the inner dimensionality of the LLM at hand.
While kadlčík2025pretrainedlanguagemodelslearn demonstrate that such a probe properly benefits from the expected sinusoidal inductive bias and greatly bolster accuracy for number representations, their approach does not enforce a strict sinusoidal structure and may fail to generalize due to redundancy in a model’s representation. As an alternative, we design a new probe which assigns a single sine wave to each latent feature and directly optimize the frequency, phase-shift, and amplitude of each sine wave:
$f_{\mathrm{param}\text{--}\sin}(\mathbf{x})$ $=\mathbf{S}^{T}(\mathbf{W}_{\mathrm{in}}\mathbf{x})$ (2)
$\mathbf{S}$ $=\mathbf{w}_{a}\sin(\mathbf{w}_{\phi}+\mathbf{w}_{f}\mathbf{r}^{T})$
where $\mathbf{r}$ are constant label values (09990...999 for all models we analyze), and $\mathbf{w}_{f},\mathbf{w}_{a},\mathbf{w}_{\phi}\in\mathbb{R}^{d}$ are three vectors of learnable parameters modeling the frequencies, amplitudes and the phase-shifts of the sine waves. This architecture strictly constrains the classification matrix $\mathbf{S}$ to take the form of $d$ sine waves sampled at $n$ evenly-spaced points. The configuration of initialization and hyperparameters is provided in Appendix B.
@@
We first assess whether sinusoidal probes are the most suitable, i.e., an accurate choice for decoding the internal representations of the model. To contextualize the accuracy in terms of sinusoidal quality of the representation, we also evaluate other types of probes used to decode numbers in previous work Feng et al. (2023); Zhu et al. (2025); kadlčík2025pretrainedlanguagemodelslearn as baselines. We fit separate probes for activations from every layer across six different language models from three families. Contrary to previous work, we assess the representations of numbers in natural-language contexts, rather than as single tokens or synthetic arithmetical commands. Our contexts cover four domains (temporal, arithmetic, medical and culinary) covering numerical contexts including quantities, measurements or timestamps (detailed in Appendix A.2). To assess generalization rather than memorization capacity, we split the prompts and corresponding representations into training, validation and test sets containing distinct extracted numbers ($x_{2}$).
@@
Figure 3: Accuracy of decoding numbers from internal activations of language models
LABEL:fig:per_layer_acc_across_probes shows the accuracy of different probes extracting the value of input numbers from representations of each model’s layer, evaluated across six different models. As we can see, both sinusoidal probes reach over 80% of accuracy in all but two cases and over 90% of accuracy for a majority of probing scenarios. Here, the superior accuracy of both sinusoidal probes provides an indicator of both (i) the sinusoidal character, and (ii) the accuracy of input number representation across layers — the best probe results present a lower bar of the accuracy of numeric representation on a given layer.
LABEL:fig:per_layer_acc_across_models_natural disentangles the accuracies of the param-sin probe from LABEL:fig:per_layer_acc_across_probes by the model, covering six LLMs of different sizes and three different families. While the probing is slightly less accurate on some particular layers of particular models, for none of the analyzed models does the sinusoidal probe systematically underperform. If we take into account that our probes need not be perfect and reflect a lower bound of real accuracy of numeric representations, these results taken together show that a sinusoidal representation is a decisively dominant scheme of numeric representations in language models, corroborating and extending the observations of Kantamneni and Tegmark (2025).
A vast majority of cases performing lower than 80% occurs in probing the models’ first and a last layer; in Section˜A.6, we analyze these cases in detail through the lens of probes learned representations and find the substantial qualitative difference in the sparsity of representations of numbers. While the middle layers of all models resort to a ‘dense’ representation decomposed into a small set of features bearing full numeric information, the initial and final layers scatter the numeric information across most of the latent features. As the scattered features share the information with natural language, a robust recovery in a realistic setting of natural contexts becomes more difficult.
Generalization across contexts
We can see that sinusoidal probes trained with natural-language contexts can provide highly accurate estimates of LLMs’ numerical states across models. However, favoring simplicity, all of the previous work (§2) investigated numerical representations in synthetic mathematical contexts. To assess whether such a simplification comes for a price in practical applicability of resulting probes, we explore whether the choice of training contexts affects generalization of resulting probes.
@@
Figure 4: Generalization of probes fitted on representations in natural language (yellow), and synthetic, mathematical contexts (blue), evaluated on natural-language (solid) and math contexts (dashed). Llama 3.2 1B.
@@
In Figure˜4, we compare the performance of probes trained on natural-language contexts (yellow) and synthetically-constructed math contexts (blue), evaluated on natural-language inputs (solid lines) and synthetic inputs (dashed lines). We can see that probes fitted with natural-language contexts are much more robust in application across both math and natural-language context. While both probes generalize well within their training context type, only the natural-language probes generalize across evaluation settings, robust to the context’s distribution shift. This observation has two important implications. First, it informs future work aiming to explain models’ internal computations in practical applications to favor probes trained in natural-language settings. Second, it restrains future work in interpretability and explainability from drawing general conclusions on models’ mechanics from fully-synthetic settings, evidencing that such conclusions may not generalize to real-world applications.
@@
In summary, our results underscore that the same type of sinusoidal representation of numbers holds in general across different model types. Numbers are represented in a similar, systematic fashion, regardless of which layer or type of context we consider. While we can expect this to be in part due to the residual stream across layers, we note that there is no a priori reason for an LLM to preserve sinusoidality across layers, and previous work shows that input tokens tend to have a negligible impact on upper layer representations (Mickus et al., 2022).